016.66 points 1 Previous Answers HRW10 4.P003. A positron undergoes a displacement Δ. 8.0i-45 ^ + 4 ok, ending with the position vector r-8.5i 7 ok n meters what was the positrons on er posit on vecto The displacement vector is equal to the subtraction of the initial position vector from the final position vector. You want the initial position vector Vectors are to be combie axis alone. Then combine y components alone. Finally, combine z components alone. Additional Materials EA Section 4.1 H Algebra Squareroots & Squares
Answer #1

Since ΔR=R2−R1, we may let the initial position vector be →R=xˆi+yˆj+zˆkand then compare terms to write

x−0=8⇒x=8

8.5−y=−4.5⇒y=13

−7−z=4⇒z=−11

R1=8i+13j−11k